If interest in the financial market on the Ivoirian Treasury bills is 12 percent and the expected inflation is 5 percent, what is your real rate of interest? 
Answer #1

The real rate of interest is given by:

1+ Rreal = (1+Rinterest)/(1+R inflation)

1+Rreal = (1+12%)/(1+5%)

Rreal = 1.0667 - 1 = 0.0667

Rreal = 6.67%
Answer #2

SOLUTION :


Real rate 

= (1 + Treasury bills rate) / ( 1 + inflation rate) - 1

= (1 + 0.12) / (1 + 0.05) - 1

= 0.0667

= 6.67% (ANSWER).  

answered by: Tulsiram Garg